OpenStudy (anonymous):

If the perimeter of the triangle is equal to the perimeter of the square then the length of a side of the square is ?

OpenStudy (anonymous):

HELP!!!!!!!!!!!!

OpenStudy (mathstudent55):

Assuming a regular triangle. Triangle: PT = 3S (S = Side of triangle) Square: PS = 4s (s = side of squaere) PT = PS (perimeter of triangle = perimeter of square), so 4s = 3S Divide both sides by 4, s = (3/4)S

OpenStudy (anonymous):

A. 7 B. 13 C.26 D.. 52

OpenStudy (mathstudent55):

Was there a drawing with a triangle and a square and the lengths of something given? Without knowing any specific information, there isn't any more we can do.

OpenStudy (anonymous):

yes there was

OpenStudy (anonymous):

do anyone know the answer

OpenStudy (mathstudent55):

How can anyone know the answer? This is like asking the following question: One segment is half the length of another segement. What is the length of the shorter segment?

OpenStudy (anonymous):

ok one side of the triangle is 16 the other 4x - 3 then 3x the square 2x then 2x

OpenStudy (mathstudent55):

Ok, then forget what we wrote above about the sides of the traingle and the square. The perimeter of the triangle is 16 + 4x - 3 + 3x The perimeter of the square is 4(2x) Since the two perimeters are equal, then 16 + 4x - 3 + 3x = 4(2x) Now you need to solve for x. The the length of the side of the square is 2x

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

16 + 4x - 3 + 3x = 4(2x) 13 + 7x = 8x 13 = x x = 13, so 2x = 26, which is the length of the side of the square.

OpenStudy (anonymous):

ok thx